# UniCUE: Unified Recognition and Generation Framework for Chinese Cued Speech Video-to-Speech Generation

### v3

@@ Abstract.
Cued Speech (CS) enhances lipreading via hand coding, offering visual phonemic cues that support precise speech perception for the hearing-impaired. The task of CS Video-to-Speech generation (CSV2S) aims to convert CS videos into intelligible speech signals. Most existing research focuses on CS Recognition (CSR), which transcribes video content into text. Consequently, a common solution for CSV2S is to integrate CSR with a text-to-speech (TTS) system. However, this pipeline relies on text as an intermediate medium, which may lead to error propagation and temporal misalignment between speech and CS video dynamics. In contrast, directly generating audio speech from CS video (direct CSV2S) often suffer from the inherent multimodal complexity and the limited availability of CS data. To address these challenges, we propose UniCUE, the first unified framework for CSV2S that directly generates speech from CS videos without relying on intermediate text. The core innovation of UniCUE lies in integrating a understanding task (CSR) that provides fine-grained CS visual-semantic cues to to guide the speech generation. Specifically, UniCUE incorporates a pose-aware visual processor, a semantic alignment pool that enables precise visual–semantic mapping, and a VisioPhonetic adapter to bridge the understanding and generation tasks within a unified architecture. To support this framework, we construct UniCUE-HI, a large-scale Mandarin CS dataset containing 11,282 videos from 14 cuers, including both hearing-impaired and normal-hearing individuals. Extensive experiments conducted on this dataset demonstrate that UniCUE achieves state-of-the-art (SOTA) performance across multiple evaluation metrics.
Chinese Cued Speech, Unified Framework, Video-to-Speech Generation, Understanding and Generation, Cued Speech Dataset
@@
Figure 1. Illustration of the rules of the Chinese CS system and the proposed framework (UniCUE). (a) The chart for Mandarin Chinese CS (figure from (Liu and Feng, 2019)), where five distinct hand positions are used to encode vowels, and eight finger shapes are employed to represent consonants in Mandarin Chinese. (b) Our framework enables the direct generation of synchronized natural speech from video.
1. Introduction
@@
Cued Speech (CS) is an visual phonetic encoding system that utilizes specific hand shapes and positions to enhance lip reading, providing an accurate visual representation of all phonemes in spoken language (Cornett, 1967; Liu and Feng, 2019; Leybaert and LaSasso, 2010). CS maintains a high level of consistency with spoken language in terms of phonemes and speech patterns, enabling hearing-impaired individuals to better integrate into speech-dominant social and educational environments (Cornett, 1967; Leybaert and LaSasso, 2010; Leybaert et al., 2010). In Mandarin Chinese, CS employs 8 hand shapes and 5 positions to encode consonants and vowels (as illustrated in Figure 1(a)), addressing challenges such as the phonemes with similar lip shapes (Liu and Feng, 2019).
@@
CS Video-to-Speech generation (CSV2S) task aims to convert CS videos of into comprehensible speech signals. However, directly constructing an end-to-end CSV2S model faces several challenges. Firstly, this task involves complex multimodal semantic correlations, requiring precise mapping from visual cues (lip movements and hand coding) to acoustic speech, while the limited scale of existing CS datasets further constrains model capacity. Secondly, fine-grained spatiotemporal modeling of visual information is essential to resolve the intrinsic asynchrony, i.e., the hand-preceding phenomenon, where hand cues precede corresponding lip movements (Liu et al., 2020a). To the best of our knowledge, the CSV2S task has not been explicitly studied in prior literature.
@@
Existing research primarily focuses on CS Recognition (CSR) that converts CS videos into phoneme-level text (Liu et al., 2020a, 2024, 2018, 2019), neglecting the critical need for natural speech generation. This limitation significantly impairs real-time communication between hearing-impaired and normal-hearing individuals, especially in educational and social scenarios. For instance, in group conversations, normal-hearing participants must quickly comprehend and respond to questions posed by their hearing-impaired peers. Textual output from CSR systems is often insufficient for such natural and smooth interactions. Additionally, recent lipreading-based video-to-speech models such as LipVoicer (Yemini et al., 2024) rely solely on lip movements, failing to capture the complementary hand-coded information in CS that conveys critical phonemic distinctions. These shortcomings underscore the need for a more comprehensive approach. Motivated by this, we aim to develop the first Chinese CSV2S system that directly decodes CS videos into intelligible speech, as illustrated in Figure 1(b).
@@
Figure 2. (a) The combined CSV2S architecture combines separately trained CSR and TTS models. (b) Our unified framework (UniCUE) that transfers understanding capabilities of CSR into speech generation training by integrating the visual processor of CSR into CSV2S.
@@
To overcome these challenges, we draw inspiration from recent advances in multimodal learning, where semantic reasoning from vision-language models (VLMs) has shown strong promise in tasks like text-guided image synthesis with interleaved control (Mi et al., 2025; Chen et al., 2025a). We hypothesize that the multimodal visual understanding inherent in CSR can serve as a semantic bridge to support more accurate and controllable speech generation in CSV2S. As depicted in Figure 2(b), we introduce a unified framework that leverages a shared visual processor to bridge CSR (understanding task) and CSV2S (generation task). This processor serves as a two-way translator: during CSR, it extracts linguistic semantics from fine-grained lip-hand motion patterns; in CSV2S, it utilizes these semantics to guide speech generation. The core innovation of our framework lies in modeling a semantic compensation flow, where phoneme-level supervision from CSR reduces ambiguity in speech synthesis, enabling more faithful and coherent voice generation under complex multimodal conditions.
@@
Building upon this semantic compensation paradigm, in this work, we propose UniCUE, the first unified framework that bridges CSR and CSV2S tasks through three specific components: Firstly, unlike prior CSR methods (Liu et al., 2024; Liu and Liu, 2023) that process lip and hand modalities independently and rely on raw video embeddings, UniCUE employs a pose-aware processor that fuses video and pose streams into a mixed representation. This enables fine-grained spatiotemporal modeling of the hand-preceding phenomenon and improves generalization to cuer-specific expressive styles. Secondly, to enhance the alignment between visual and linguistic semantics, we introduce a semantic alignment pool to map the video and pose latent spaces into a shared textual space using contrastive learning. This facilitates cross-modal correlation modeling and improves semantic consistency in the generated speech. Thirdly, to unify the understanding and generation tasks, we reuse the CSR visual encoder within our diffusion-based CSV2S decoder and introduce a VisioPhonetic Adapter (VPA) that transforms the visual representations into diffusion-compatible codes. This design enables the decoder to effectively incorporate fine-grained semantic information derived from multimodal visual inputs
@@
To evaluate UniCUE on hearing-impaired individuals, we extend the MCCS dataset (Lei et al., 2024b) by adding data from 8 hearing-impaired and 2 normal-hearing cuers [^1] Cuer means the people who perform CS., forming the Unified-HI Corpus with 14 cuers. Experimental results on this dataset demonstrate that UniCUE not only produces accurate and intelligible speech, but also maintains temporal synchronization with the CS video.
@@
Figure 3. Overview of our unified framework (UniCUE). It achieves direct Chinese CSV2S generation with semantic consistency, temporal alignment, and characteristics coherence by aligning the fine-grained spatiotemporal visual representations of CSR with the diffusion-based speech generator. The framework consists of three core modules: (1) Pose-Aware Visual Processor: Integrates video and pose embeddings to perform fine-grained spatiotemporal modeling of lip and hand movements. (2) Semantic Alignment Pool: Enhances the semantic mapping between visual features and speech content through video-text and pose-text contrastive learning. (3) VisioPhonetic Adapter (VPA): Converts fine-grained visual representation of CSR into condition encodings compatible with the diffusion-based generator.
2. Related Work
2.1. Video-to-Speech Generation
@@
V2S aims to synthesize natural speech aligned with silent talking videos, but is challenged by limited data. Uni-Dubbing (Lei et al., 2024a) addresses this via modality-aligned pre-training on multimodal data and fine-tuning with both multimodal and audio-only inputs. Similarly, Kefalas et al. (Kefalas et al., 2024) pre-train on large audio-only corpora before tuning on paired data. Some studies (Kim et al., 2023; Yemini et al., 2024; Gupta et al., 2024) incorporate transcripts to enhance generation. Kim et al. (Kim et al., 2023) use text-speech supervision to improve word-level representation via multi-task learning. Existing V2S methods primarily focus on lip reading. However, CS conveys phonemic information through both lip and hand movements. Ignoring hand cues results in incomplete visual representations and degraded speech synthesis quality, limiting the applicability of these methods to CS. Notably, no prior work has addressed the CSV2S task.
2.2. Cued Speech Recognition
@@
CS augments lip reading with hand coding to support the hearing-impaired. The CSR task aims to transcribe CS videos into text by leveraging lips and hands as complementary modalities (Liu and Feng, 2019; Papadimitriou and Potamianos, 2021). Most CSR methods extract lip and hand features separately and fuse them for recognition (Liu et al., 2020a, 2024; Liu and Liu, 2023; Zhang et al., 2023). Due to the asynchronous nature of these modalities, effective fusion remains challenging. Liu et al. (Liu et al., 2020a) proposed re-synchronization to align hand with lip features, while transformer-based mutual learning (Liu and Liu, 2023; Liu et al., 2024) improves multimodal interaction. Zhang et al. (Zhang et al., 2023) addressed privacy concerns via federated learning. In contrast, we directly model lip and hand cues from whole frames, avoiding explicit fusion. A pose-aware visual processor is introduced further to enhance cross-modal representation and improve performance.
2.3. Unified Understanding and Generation
@@
Recent advances in unifying understanding and generation tasks fall into two main paradigms. The first integrates visual-language understanding with external generative models (e.g., diffusion models) for multimodal generation (Wu et al., 2024a; Dong et al., 2024; Jin et al., 2024; Li et al., 2024; Sun et al., 2024; Ge et al., 2024, 2025). For example, (Jin et al., 2024; Li et al., 2024) utilize large language models (LLMs) for semantic understanding and diffusion models (Rombach et al., 2022; Podell et al., 2023) for high-fidelity image synthesis. The second paradigm trains LLM-based foundation models via next-token prediction for both vision understanding and generation (Yu et al., 2023; Sun et al., 2023; Zhou et al., 2024; Wu et al., 2024b; Fang et al., 2024; Wu et al., 2025; Chen et al., 2025b). Transfusion (Zhou et al., 2024), for instance, unifies image understanding and generation within a single transformer, enabling controllable text-to-image synthesis by preserving visual details. However, existing approaches mainly focus on visual-text settings, leaving visual-to-speech generation underexplored. In this work, we introduce the first unified framework that bridges visual understanding and speech generation.
3. Method
3.1. Overview of UniCUE
@@
To achieve accurate CSV2S generation, the proposed method needs to simultaneously address two critical challenges: (1) semantic understanding of the linguistic correlations between visual cues and speech content, and (2) speech synthesis that preserves cuer-specific characteristics and temporal alignment. Inspired by the auxiliary benefits of unified understanding and generation for multi-modal controllable image synthesis (Wu et al., 2024a; Dong et al., 2024), we design a unified architecture that integrates CSR and CSV2S, enabling CSV2S with understanding capability improvement through shared visual feature representations. As illustrated in Figure 3, the framework operates via two pathways.
@@
CSR: Fine-grained Visual Cues Understanding. As the recognition pathway, CSR models fine-grained spatiotemporal visual semantics to transcribe CS videos into linguistic sequences. Given a CS video $I_{v}$ and its corresponding pose maps $I_{p}$ (extracted via OpenPose (Cao et al., 2019)), we first utilize a pose-aware visual processor to extract multi-modal embeddings $Z_{mv}$, which capture lip and hand motion cues. And then $Z_{mv}$ is fed into a auto-regressive Transformer-based text decoder $D_{T}$, which models long-range dependencies and contextual interactions across the sequence to generate the predicted token sequence: $T_{p}=D_{T}(Z_{mv})$, where $T_{p}$ denotes the predicted token sequence.
@@
Unlike prior approaches relying on Connectionist Temporal Classification (CTC) loss (Graves et al., 2006), which predict each token independently and thus limit the model’s ability to capture cross-token dependencies and coarticulatory effects, our method employs an auto-regressive decoder $D_{T}$ supervised by cross-entropy loss. This design allows $D_{T}$ to generate tokens conditioned on previously generated outputs and spatialtemporal visual cues, which is more suited to modeling the asynchronous and dynamic nature of CS.
@@
(1) $\mathcal{L}_{R}=\mathcal{L}_{\text{CE}}^{\text{masked}}(T_{p},T_{g})+\mathcal{L}_{\text{CE}}^{\text{seq}}(T_{p},T_{g}),$
@@
CSV2S: Cuer-specific Speech Synthesis. To directly synthesize intelligible and personalized speech from CS videos, we formulate speech generation as a conditional denoising process within a latent diffusion model (LDM) (Rombach et al., 2022). Since both lip shapes and hand cues in CS convey phonemic content, the speech generation is conditioned a refined visual embedding $Z_{mv}^{\prime}$, which is derived by transforming the CSR multimodal feature $Z_{mv}$ via a VisioPhonetic adapter (VPA). Specifically, a pretrained VAE encoder compresses ground-truth mel-spectrograms into latent codes $Z_{s}$, which are progressively corrupted with Gaussian noise $\epsilon$ over $t$ steps: $Z_{s}^{t}:=\alpha_{t}\cdot Z_{s}+(1-\alpha_{t})\cdot\epsilon,$ where $\alpha_{t}$ denotes the noise level at timestep $t$. The noisy latent $Z_{s}^{t}$ then denoised by the LDM conditioned on $Z_{mv}^{\prime}$. The generation objective is defined as:
(2) $\mathcal{L}_{G}:=\mathbb{E}_{Z_{s}^{t},\,Z_{mv},\,\epsilon,\,t}\left[\left\|\epsilon-\mathcal{M}(Z_{s}^{t},Z_{mv},t)\right\|^{2}_{2}\right],$
@@
3.2. Pose-aware Visual Processor
@@
(3) $Z_{mv}=\text{MLP}(\text{Concat}(Z_{v},Z_{p})).$
@@
3.3. Semantic Alignment Pool
@@
(4) $\mathcal{L}_{v\leftrightarrow t}=1-\cos(Z_{v},Z_{t}),\quad\mathcal{L}_{p\leftrightarrow t}=1-\cos(Z_{p},Z_{t}),$
@@
(5) $\mathcal{L}_{S}=\mathcal{L}_{v\leftrightarrow t}+\mathcal{L}_{p\leftrightarrow t}.$
@@
Figure 4. The details of the VisioPhonetic Adapter, which transforms semantic visual embeddings into phonetic-aware features to enable seamless conditioning for diffusion-based speech synthesis.
3.4. VisioPhonetic Adapter
@@
(6) $\mathbf{Z}_{mv}^{\prime}=\text{MLP}\Big{(}\text{CrossAttn}\big{(}\text{MLP}(\mathbf{Z}_{mv})\big{)}\Big{)},$
which includes two MLPs and a Q-Former-style (Li et al., 2023) cross-attention layer. We use $N_{q}$ learnable semantic queries $\mathbf{f}\in\mathbb{R}^{N_{q}\times D}$, which is initialized by computing the average latent representation from ground-truth mel-spectrograms encoded by the pretrained VAE. This provides a phonetic-aware initialization aligned with the diffusion model’s target space. These queries act as phonetic slots to extract and reorganize relevant patterns from $\mathbf{Z}_{mv}$. The cross-attention mechanism operates as: $\mathbf{q}=\mathbf{W}^{q}\mathbf{f},\mathbf{k}=\mathbf{W}^{k}\mathbf{Z_{mv}},\mathbf{v}=\mathbf{W}^{v}\mathbf{Z_{mv}},\mathbf{a}=\text{Softmax}\left(\frac{\mathbf{q}\mathbf{k}^{T}}{\sqrt{d}}\right)\mathbf{v},\mathbf{{Z_{mv}}^{\prime}}=\text{MLP}(\mathbf{Z_{mv}}+a).$ The adapted features $\mathbf{Z}_{mv}^{\prime}$ serve as the final interface between visual understanding and speech synthesis, ensuring that the generated audio is not only temporally coherent but also linguistically faithful to the video input.
Table 1. Comparison between our Chinese Mandarin CS dataset and existing CS dataset. H denotes the cuers with normal hearing, while HI indicates hearing-impaired cuer. Our newly proposed Unified-HI Corpus is the first large-scale Chinese CS dataset with both hearing-impaired and normal-hearing cuers.
@@
  French CS (Liu et al., 2018) 1-H 238 12872 - $720\times 576$ 50
British CS (Liu et al., 2019) 1-H 97 2741 - $720\times 1280$ 25
MCCS (Lei et al., 2024b) 4-H 4000 131608 42256 $720\times 1280$ 30
@@
Table 2. Comparison with SOTA methods on test data of normal-hearing cuers and hearing-impaired cuers. Bold and underlined results are the best and second-best results. $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are preferable.
@@
4. Experiment
4.1. Experimental Setting

### v4

@@ Abstract
Cued Speech (CS) enhances lipreading via hand coding, offering visual phonemic cues that support precise speech perception for the hearing-impaired. The task of CS Video-to-Speech generation (CSV2S) aims to convert CS videos into intelligible speech signals. Most existing research focuses on CS Recognition (CSR), which transcribes video content into text. Consequently, a common solution for CSV2S is to integrate CSR with a text-to-speech (TTS) system. However, this pipeline relies on text as an intermediate medium, which may lead to error propagation and temporal misalignment between speech and CS video dynamics. In contrast, directly generating audio speech from CS video (direct CSV2S) often suffer from the inherent multimodal complexity and the limited availability of CS data. To address these challenges, we propose UniCUE, the first unified framework for CSV2S that directly generates speech from CS videos without relying on intermediate text. The core innovation of UniCUE lies in integrating a understanding task (CSR) that provides fine-grained CS visual-semantic cues to to guide the speech generation. Specifically, UniCUE incorporates a pose-aware visual processor, a semantic alignment pool that enables precise visual–semantic mapping, and a VisioPhonetic adapter to bridge the understanding and generation tasks within a unified architecture. To support this framework, we construct UniCUE-HI, a large-scale Mandarin CS dataset containing 11,282 videos from 14 cuers, including both hearing-impaired and normal-hearing individuals. Extensive experiments conducted on this dataset demonstrate that UniCUE achieves state-of-the-art (SOTA) performance across multiple evaluation metrics. Project website can be found at https://beria-moon.github.io/UniCUE/
@@
Figure 1: Illustration of the rules of the Chinese CS system and the proposed framework (UniCUE). (a) The chart for Mandarin Chinese CS (figure from (liu2019pilot)), where five distinct hand positions are used to encode vowels, and eight finger shapes are employed to represent consonants in Mandarin Chinese. (b) Our framework enables the direct generation of synchronized natural speech from video.
Introduction
@@
Cued Speech (CS) is an visual phonetic encoding system that utilizes specific hand shapes and positions to enhance lip reading, providing an accurate visual representation of all phonemes in spoken language (cornett1967cued; liu2019pilot; leybaert2010cued). CS maintains a high level of consistency with spoken language in terms of phonemes and speech patterns, enabling hearing-impaired individuals to better integrate into speech-dominant social and educational environments (cornett1967cued; leybaert2010cued; leybaert201019). In Mandarin Chinese, CS employs 8 hand shapes and 5 positions to encode consonants and vowels (as illustrated in Figure 1(a)), addressing challenges such as the phonemes with similar lip shapes (liu2019pilot).
@@
CS Video-to-Speech generation (CSV2S) task aims to convert CS videos of into comprehensible speech signals. However, directly constructing an end-to-end CSV2S model faces several challenges. Firstly, this task involves complex multimodal semantic correlations, requiring precise mapping from visual cues (lip movements and hand coding) to acoustic speech, while the limited scale of existing CS datasets further constrains model capacity. Secondly, fine-grained spatiotemporal modeling of visual information is essential to resolve the intrinsic asynchrony, i.e., the hand-preceding phenomenon, where hand cues precede corresponding lip movements (liu2020re). To the best of our knowledge, the CSV2S task has not been explicitly studied in prior literature.
@@
Existing research primarily focuses on CS Recognition (CSR) that converts CS videos into phoneme-level text (liu2020re; liu2024computation; liu2018visual; liu2019automatic), neglecting the critical need for natural speech generation. This limitation significantly impairs real-time communication between hearing-impaired and normal-hearing individuals, especially in educational and social scenarios. For instance, in group conversations, normal-hearing participants must quickly comprehend and respond to questions posed by their hearing-impaired peers. Textual output from CSR systems is often insufficient for such natural and smooth interactions. Additionally, recent lipreading-based video-to-speech models such as LipVoicer (yeminilipvoicer) rely solely on lip movements, failing to capture the complementary hand-coded information in CS that conveys critical phonemic distinctions. These shortcomings underscore the need for a more comprehensive approach. Motivated by this, we aim to develop the first Chinese CSV2S system that directly decodes CS videos into intelligible speech, as illustrated in Figure 1(b).
@@
Figure 2: (a) The combined CSV2S architecture combines separately trained CSR and TTS models. (b) Our unified framework (UniCUE) that transfers understanding capabilities of CSR into speech generation training by integrating the visual processor of CSR into CSV2S.
@@
To overcome these challenges, we draw inspiration from recent advances in multimodal learning, where semantic reasoning from vision-language models (VLMs) has shown strong promise in tasks like text-guided image synthesis with interleaved control (mi2025think; chen2025multimodal). We hypothesize that the multimodal visual understanding inherent in CSR can serve as a semantic bridge to support more accurate and controllable speech generation in CSV2S. As depicted in Figure 2(b), we introduce a unified framework that leverages a shared visual processor to bridge CSR (understanding task) and CSV2S (generation task). This processor serves as a two-way translator: during CSR, it extracts linguistic semantics from fine-grained lip-hand motion patterns; in CSV2S, it utilizes these semantics to guide speech generation. The core innovation of our framework lies in modeling a semantic compensation flow, where phoneme-level supervision from CSR reduces ambiguity in speech synthesis, enabling more faithful and coherent voice generation under complex multimodal conditions.
@@
Building upon this semantic compensation paradigm, in this work, we propose UniCUE, the first unified framework that bridges CSR and CSV2S tasks through three specific components: Firstly, unlike prior CSR methods (liu2024computation; liu2023cross) that process lip and hand modalities independently and rely on raw video embeddings, UniCUE employs a pose-aware processor that fuses video and pose streams into a mixed representation. This enables fine-grained spatiotemporal modeling of the hand-preceding phenomenon and improves generalization to cuer-specific expressive styles. Secondly, to enhance the alignment between visual and linguistic semantics, we introduce a semantic alignment pool to map the video and pose latent spaces into a shared textual space using contrastive learning. This facilitates cross-modal correlation modeling and improves semantic consistency in the generated speech. Thirdly, to unify the understanding and generation tasks, we reuse the CSR visual encoder within our diffusion-based CSV2S decoder and introduce a VisioPhonetic Adapter (VPA) that transforms the visual representations into diffusion-compatible codes. This design enables the decoder to effectively incorporate fine-grained semantic information derived from multimodal visual inputs
@@
To evaluate UniCUE on hearing-impaired individuals, we extend the MCCS dataset (lei2024bridge) by adding data from 8 hearing-impaired and 2 normal-hearing cuers [^2] Cuer means the people who perform CS., forming the Unified-HI Corpus with 14 cuers.
Experimental results on this dataset demonstrate that UniCUE not only produces accurate and intelligible speech, but also maintains temporal synchronization with the CS video.
@@
Figure 3: Overview of our unified framework (UniCUE). It achieves direct Chinese CSV2S generation with semantic consistency, temporal alignment, and characteristics coherence by aligning the fine-grained spatiotemporal visual representations of CSR with the diffusion-based speech generator. The framework consists of three core modules: (1) Pose-Aware Visual Processor: Integrates video and pose embeddings to perform fine-grained spatiotemporal modeling of lip and hand movements. (2) Semantic Alignment Pool: Enhances the semantic mapping between visual features and speech content through video-text and pose-text contrastive learning. (3) VisioPhonetic Adapter (VPA): Converts fine-grained visual representation of CSR into condition encodings compatible with the diffusion-based generator.
Related Work
Video-to-Speech Generation
@@
V2S aims to synthesize natural speech aligned with silent talking videos, but is challenged by limited data. Uni-Dubbing (lei2024uni) addresses this via modality-aligned pre-training on multimodal data and fine-tuning with both multimodal and audio-only inputs. Similarly, Kefalas et al. (kefalas2024large) pre-train on large audio-only corpora before tuning on paired data. Some studies (kim2023lip; yeminilipvoicer; gupta2024visatronic) incorporate transcripts to enhance generation. Kim et al. (kim2023lip) use text-speech supervision to improve word-level representation via multi-task learning. Existing V2S methods primarily focus on lip reading. However, CS conveys phonemic information through both lip and hand movements. Ignoring hand cues results in incomplete visual representations and degraded speech synthesis quality, limiting the applicability of these methods to CS. Notably, no prior work has addressed the CSV2S task.
Cued Speech Recognition
@@
CS augments lip reading with hand coding to support the hearing-impaired. The CSR task aims to transcribe CS videos into text by leveraging lips and hands as complementary modalities (liu2019pilot; papadimitriou2021fully). Most CSR methods extract lip and hand features separately and fuse them for recognition (liu2020re; liu2024computation; liu2023cross; zhang2023cuing). Due to the asynchronous nature of these modalities, effective fusion remains challenging. Liu et al. (liu2020re) proposed re-synchronization to align hand with lip features, while transformer-based mutual learning (liu2023cross; liu2024computation) improves multimodal interaction. Zhang et al. (zhang2023cuing) addressed privacy concerns via federated learning. In contrast, we directly model lip and hand cues from whole frames, avoiding explicit fusion. A pose-aware visual processor is introduced further to enhance cross-modal representation and improve performance.
Unified Understanding and Generation
@@
Recent advances in unifying understanding and generation tasks fall into two main paradigms. The first integrates visual-language understanding with external generative models (e.g., diffusion models) for multimodal generation (wu2024next; dong2024dreamllm; jin2024unified; li2024mini; sun2024generative; ge2024seed; ge2025divot). For example, (jin2024unified; li2024mini) utilize large language models (LLMs) for semantic understanding and diffusion models (rombach2022high; podell2023sdxl) for high-fidelity image synthesis. The second paradigm trains LLM-based foundation models via next-token prediction for both vision understanding and generation (yu2023scaling; sun2023emu; zhoutransfusion; wu2024vila; fang2024puma; wu2025janus; chen2025janus). Transfusion (zhoutransfusion), for instance, unifies image understanding and generation within a single transformer, enabling controllable text-to-image synthesis by preserving visual details. However, existing approaches mainly focus on visual-text settings, leaving visual-to-speech generation underexplored. In this work, we introduce the first unified framework that bridges visual understanding and speech generation.
Method
Overview of UniCUE
@@
To achieve accurate CSV2S generation, the proposed method needs to simultaneously address two critical challenges: (1) semantic understanding of the linguistic correlations between visual cues and speech content, and (2) speech synthesis that preserves cuer-specific characteristics and temporal alignment. Inspired by the auxiliary benefits of unified understanding and generation for multi-modal controllable image synthesis (wu2024next; dong2024dreamllm), we design a unified architecture that integrates CSR and CSV2S, enabling CSV2S with understanding capability improvement through shared visual feature representations. As illustrated in Figure 3, the framework operates via two pathways.
@@
CSR: Fine-grained Visual Cues Understanding. As the recognition pathway, CSR models fine-grained spatiotemporal visual semantics to transcribe CS videos into linguistic sequences. Given a CS video $I_{v}$ and its corresponding pose maps $I_{p}$ (extracted via OpenPose (8765346)), we first utilize a pose-aware visual processor to extract multi-modal embeddings $Z_{mv}$, which capture lip and hand motion cues. And then $Z_{mv}$ is fed into a auto-regressive Transformer-based text decoder $D_{T}$, which models long-range dependencies and contextual interactions across the sequence to generate the predicted token sequence: $T_{p}=D_{T}(Z_{mv})$, where $T_{p}$ denotes the predicted token sequence.
@@
Unlike prior approaches relying on Connectionist Temporal Classification (CTC) loss (graves2006ctc), which predict each token independently and thus limit the model’s ability to capture cross-token dependencies and coarticulatory effects, our method employs an auto-regressive decoder $D_{T}$ supervised by cross-entropy loss. This design allows $D_{T}$ to generate tokens conditioned on previously generated outputs and spatialtemporal visual cues, which is more suited to modeling the asynchronous and dynamic nature of CS.
@@
$\mathcal{L}_{R}=\mathcal{L}_{\text{CE}}^{\text{masked}}(T_{p},T_{g})+\mathcal{L}_{\text{CE}}^{\text{seq}}(T_{p},T_{g}),$ (1)
@@
CSV2S: Cuer-specific Speech Synthesis. To directly synthesize intelligible and personalized speech from CS videos, we formulate speech generation as a conditional denoising process within a latent diffusion model (LDM) (rombach2022high). Since both lip shapes and hand cues in CS convey phonemic content, the speech generation is conditioned a refined visual embedding $Z_{mv}^{\prime}$, which is derived by transforming the CSR multimodal feature $Z_{mv}$ via a VisioPhonetic adapter (VPA). Specifically, a pretrained VAE encoder compresses ground-truth mel-spectrograms into latent codes $Z_{s}$, which are progressively corrupted with Gaussian noise $\epsilon$ over $t$ steps: $Z_{s}^{t}:=\alpha_{t}\cdot Z_{s}+(1-\alpha_{t})\cdot\epsilon,$ where $\alpha_{t}$ denotes the noise level at timestep $t$. The noisy latent $Z_{s}^{t}$ then denoised by the LDM conditioned on $Z_{mv}^{\prime}$. The generation objective is defined as:
$\mathcal{L}_{G}:=\mathbb{E}_{Z_{s}^{t},\,Z_{mv},\,\epsilon,\,t}\left[\left\|\epsilon-\mathcal{M}(Z_{s}^{t},Z_{mv},t)\right\|^{2}_{2}\right],$ (2)
@@
Pose-aware Visual Processor
@@
$Z_{mv}=\text{MLP}(\text{Concat}(Z_{v},Z_{p})).$ (3)
@@
Semantic Alignment Pool
@@
$\mathcal{L}_{v\leftrightarrow t}=1-\cos(Z_{v},Z_{t}),\quad\mathcal{L}_{p\leftrightarrow t}=1-\cos(Z_{p},Z_{t}),$ (4)
@@
$\mathcal{L}_{S}=\mathcal{L}_{v\leftrightarrow t}+\mathcal{L}_{p\leftrightarrow t}.$ (5)
@@
Figure 4: The details of the VisioPhonetic Adapter, which transforms semantic visual embeddings into phonetic-aware features to enable seamless conditioning for diffusion-based speech synthesis.
VisioPhonetic Adapter
@@
$\mathbf{Z}_{mv}^{\prime}=\text{MLP}\Big(\text{CrossAttn}\big(\text{MLP}(\mathbf{Z}_{mv})\big)\Big),$ (6)
which includes two MLPs and a Q-Former-style (li2023blip) cross-attention layer. We use $N_{q}$ learnable semantic queries $\mathbf{f}\in\mathbb{R}^{N_{q}\times D}$, which is initialized by computing the average latent representation from ground-truth mel-spectrograms encoded by the pretrained VAE. This provides a phonetic-aware initialization aligned with the diffusion model’s target space. These queries act as phonetic slots to extract and reorganize relevant patterns from $\mathbf{Z}_{mv}$. The cross-attention mechanism operates as: $\mathbf{q}=\mathbf{W}^{q}\mathbf{f},\mathbf{k}=\mathbf{W}^{k}\mathbf{Z_{mv}},\mathbf{v}=\mathbf{W}^{v}\mathbf{Z_{mv}},\mathbf{a}=\text{Softmax}\left(\frac{\mathbf{q}\mathbf{k}^{T}}{\sqrt{d}}\right)\mathbf{v},\mathbf{{Z_{mv}}^{\prime}}=\text{MLP}(\mathbf{Z_{mv}}+a).$ The adapted features $\mathbf{Z}_{mv}^{\prime}$ serve as the final interface between visual understanding and speech synthesis, ensuring that the generated audio is not only temporally coherent but also linguistically faithful to the video input.
Table 1: Comparison between our Chinese Mandarin CS dataset and existing CS dataset. H denotes the cuers with normal hearing, while HI indicates hearing-impaired cuer. Our newly proposed Unified-HI Corpus is the first large-scale Chinese CS dataset with both hearing-impaired and normal-hearing cuers.
@@
  French CS (liu2018visual) 1-H 238 12872 - $720\times 576$ 50
British CS (liu2019automatic) 1-H 97 2741 - $720\times 1280$ 25
MCCS (lei2024bridge) 4-H 4000 131608 42256 $720\times 1280$ 30
@@
Table 2: Comparison with SOTA methods on test data of normal-hearing cuers and hearing-impaired cuers. Bold and underlined results are the best and second-best results. $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are preferable.
@@
Experiment
Experimental Setting